# Exact Simulation of Longitudinal Data from Marginal Structural Models

### v2

@@ 6 Discussion
In this paper, we proposed and demonstrated a principled strategy for simulating longitudinal data exactly from marginal structural models, leveraging the frugal parameterization (Evans and Didelez, 2024). Our proposed method is highly flexible and can accommodate a variety of practical scenarios, including multiple treatments (for example, primary and drop-in), dynamic treatment regimes, dose-response analyses, multiple outcomes (such as secondary outcomes and competing risks), and censoring. Additionally, it imposes no restrictions on the data-generating process or marginal structural causal models. In Section 5, we extended this method to survival models and time-to-event outcomes, which are common in clinical trials. Our methods are both conceptually and computationally simple, which crucial for practical applications such as sample size calculation.
[…]
Appendix A Extension to Include Time-Varying Latent Variables
We provide an example of how to incorporate a latent process that confounds the covariates and the final outcome following the algorithm outlined in Section 4. The steps to simulate longitudinal data from the model depicted in Figure 3 are as follows:
@@ Appendix A Extension to Include Time-Varying Latent Variables
Appendix B Proof of Proposition 5.1
@@
Initialize the algorithm by specifying the cumulative distribution functions $F_{L_{0}}$, $F_{A_{0}\mid L_{0}}$, $F_{\tilde{Y}_{1}\mid do(A_{0})}$, and the copula $C_{\tilde{Y}_{1}L_{0}\mid do(A_{0})}$, parameterized by $\theta_{0,0}$. Draw $w_{0,1}$, $w_{0,2}$, $w_{0,3}\sim U(0,1)$ and compute:
$l_{0}=F^{-1}_{L_{0}}(w_{0,1})$ $a_{0}=F^{-1}_{A_{0}\mid L_{0}}(w_{0,2}\,|\,l_{0}),$ (16)
@@
$F_{\tilde{Y}_{1}\,|\,do(A_{0})}(\tilde{y}_{1}\,|\,do(a_{0}))$ $=h^{-1}\big{\{}F_{\tilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\tilde{y}_{1}\,|\,l_{0},do(a_{0})),F_{L_{0}}(l_{0});\theta_{0,0}\big{\}},$ (17)
set $w_{0,3}=F_{\tilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\tilde{y}_{1}\,|\,l_{0},do(a_{0}))$ and substitute $F_{L_{0}}(l_{0})$ with $w_{0,1}$, yielding:
$u_{1}=h^{-1}\left(w_{0,3},w_{0,1};\theta_{0,0}\right).$ (18)
By inversion, compute $\tilde{y}_{1}=F_{Y_{1}\mid do(A_{0})}^{-1}(u_{1})$, and the corresponding binary outcome $y_{1}=I(\tilde{y}_{1}<1)$. This completes the first iteration, yielding a sample $(l_{0},a_{0},\tilde{y}_{1},y_{1})$. If $y_{1}=1$, stop; otherwise, proceed to $k=1$.
For $k=1$, define the necessary cumulative distribution functions $F_{L_{1}\,|\,L_{0},A_{0}}$, $F_{A_{1}\mid\bar{L}_{1},A_{0}}$, $F_{\tilde{Y}_{2}\mid do(\bar{A}_{1})}$, and copulas $C_{\tilde{Y}_{2}L_{1}\mid L_{0}do(\bar{A}_{1})}$ and $C_{\tilde{Y}_{2}L_{0}\mid do(A_{0})}$, parameterized by $\theta_{1,1}$ and $\theta_{1,0}$, respectively. These two copulas associate $\tilde{Y}_{2}$ with previous time-varying confounders. Draw $w_{1,1}$, $w_{1,2}$, $w_{1,3}$ from $U(0,1)$, and obtain $(l_{1},a_{1})$ as before.
We now need to obtain a sample of $\tilde{y}_{2}$. Calculate:
$F(l_{0}\,|\,do(a_{0}),y_{1}=0)$ $=h\big{\{}F(l_{0}),F(\tilde{y}_{1}\,|\,do(a_{0}));\theta_{0,0}\big{\}}$ (19)
$F(\tilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{2}\,|\,\bar{l}_{1},do(\bar{a}_{1}),y_{1}=0),F(l_{1}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0);\theta_{1,1}\big{\}}$
$F(\tilde{y}_{2}\,|\,do(a_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\tilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0),F(l_{0}\,|\,do(\bar{a}_{1}),y_{1}=0);\theta_{1,0}\big{\}}.$
Note that we omit some subscripts on quantiles for brevity. In (19) we derive the quantile of $L_{0}$ given $a_{0}$ among survivors the start of the $k=1$ interval, where $F(l_{0})=w_{0,0}$ and $F(\tilde{y}_{1}\,|\,do(a_{0})$ is already computed in (18) in step $k=0$. All other necessary quantities are specified at $k=1$. Therefore, the base case is satisfied since the algorithm simulates the first time step and provides the required inputs for the next.
@@
Appendix C Competing risk
At each time step $k$, after simulating time-varying covariate $L_{k}$ and treatment $A_{k}$, we generate the primary outcome $Y_{k}$ and a competing risk $D_{k}$ using their respective auxiliary variables $\tilde{Y}_{k}$ and $\tilde{D}_{k}$, following the approach outlined in Section 5. For brevity, baseline covariates $\bm{Z}$ continue to be omitted, but in settings where they are present, they can be simulated as the initial step and included in the conditioning set of all subsequent quantities.
Under this simulation framework, we specify:
  • A marginal structural model for the primary event: $F^{*}(\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
  • A marginal structural model for the competing event: $F^{*}(\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
  • A bivariate copula parameterized by $\Psi_{k+1}$ to model the dependence between the above interventional margins of $\tilde{Y}_{k+1}$ and $\tilde{D}_{k+1}$;
  • Bivariate copulas that ‘weave’ the multivariate dependence between $Y_{k+1}$ and past covariates $\bar{L}_{k}$, given $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$;
  • Bivariate copulas that ‘weave’ the multivariate dependence between $\tilde{D}_{k+1}$ and $\bar{L}_{k}$ conditional on $\tilde{Y}_{k+1}$, $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Phi_{k,j}$ for $j=0,\ldots,k$.
The joint outcome model $p(\tilde{y}_{k+1},\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$ factorizes as follows:
Primary event model:
$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$
$=p^{*}(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)\times$
$\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0,D_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\}.$ (20)
Competing event model:
$p(\tilde{d}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)$
$=p(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)\times$
$\prod_{j=0}^{k}c^{*}_{\tilde{D}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},\tilde{Y}_{k+1}D_{k}=0}\big{\{}F(d_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0);\Phi{k,j}\}.$ (21)
The first term in (21), the distribution of $\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0$ can be obtained through:
$F(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)=h\big{\{}F*(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0),F^{*}(\tilde{y}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)\big{\}}.$
The challenge lies in computing $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ in (20), and $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ in (21), for $j=0,\ldots,k$. The former can be calculated in the same iterative fashion as described in Section 5. The latter can be derived as:
$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ $=h\big{\{}F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
$\qquad F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\big{\}}.$ (22)
Notably, the two components in the $h$-function in (22) exactly correspond to the those in the primary event model in (20), except with the order of conditioning swapped. Since these values are computed during the simulation of $\tilde{y}_{k+1}$, requiring only $k+1$ additional evaluations of the $h$-functions in (22).
The algorithm for simulating each time step $k$ for an individual can be summarized into:
  1. 1.
    Simulate the time-varying confounder $L_{k}$ and treatment $A_{k}$ from their respective distributions.
  2. 2.
    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=0,\dots,k-1$.
  3. 3.
    Compute $F(y_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=1,\dots k$, leading to $u_{k+1}=F(\tilde{y}_{k+1}\,|\,z,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
  4. 4.
    Compute: $\tilde{y}_{k+1}=F^{*-1}_{\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
  5. 5.
    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ for $j=0,\ldots,k$ by evaluating $k+1$ $h$-functions following (22).
  6. 6.
    Compute $F(\tilde{d}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1}=0,D_{k}=0)$ for $j=1,\ldots k$, leading to $v_{k+1}=F(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
  7. 7.
    Compute: $\tilde{d}_{k+1}=F_{\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),\tilde{Y}_{k+1},D_{k}=0}^{*-1}(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
This approach requires at most $74k+$ 7 evaluations of analytic functions per individual per time step. An example implementation is available at https://github.com/XiLinStats/LongSiMSM.
@@ Appendix C Competing risk
Models for baseline, time-varying covariates and treatment
A baseline:

### v3

@@ 6 Discussion
In this paper, we proposed and demonstrated a principled strategy for simulating longitudinal data exactly from marginal structural models, leveraging the frugal parameterization (Evans and Didelez, 2024). Our proposed method is highly flexible and can accommodate a variety of practical scenarios, including multiple treatments (for example, primary and drop-in), dynamic treatment regimes, dose-response analyses, multiple outcomes (such as secondary outcomes and competing risks), and censoring. Additionally, it imposes no restrictions on the data-generating process or marginal structural causal models. In Section 5, we extended this method to survival models and time-to-event outcomes, which are common in clinical trials. The method’s computational efficiency, combined with its analytic precision, makes it well-suited to practical tasks such as simulation-based method evaluation and sample size determination.
In Section 5.3, we discussed how our proposed simulation framework can flexibly generate data with competing risks from marginal structural models targeting the marginal hazard. However, extending the framework to simulate data under alternative estimands—specifically, the cause-specific and subdistribution hazards, which represent total effects—remains nontrivial. Future work may explore how to extend the proposed method to support simulation from marginal structural models that target these alternative causal quantities.
A limitation of the method is difficulty in enforcing Markov-type conditional independences between the outcome and the history of time-varying confounders. For instance, the pair-copula construction we employed in Section 4.2 to model the association between $Y$ and $L_{0},\dots,L_{k}$ makes it difficult to impose restrictions that might be considered natural; for example, that the distribution of the outcome only depends upon the most recent few values of a time-varying covariate. Parameterizing such a model is infeasible unless all the bivariate copulas in the pair-copula construction are Gaussian; in this case, conditional independence can be enforced by appropriately choosing the partial correlation parameters. If enforcing such Markov conditions is absolutely necessary, we would suggest using either this Gaussian approximation or numerical integration. However, numerical integration in high-dimensions is well known to be infeasible.
Appendix A Extension to include multiple time-varying variables
We outline how the proposed algorithm in Section 4 can be extended to include multiple time-varying covariates. From a simulation perspective, there is no distinction between whether such covariate is observed or unobserved; for illustrative purposes, we assume that the additional time-varying variable $H$ follows a latent process$\{H_{k}\}_{k\in\mathcal{K}}$. The steps for simulate data from the model depicted in Figure 5 are as follows:
@@ Appendix A Extension to include multiple time-varying variables
Appendix B Proof of Lemma 5.1
Proof.
By definition:
$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k}=0)$
[…]
where $q_{k-1,j}=F(\widetilde{y}_{k}=1\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$. Recall the result from Joe (1996) in (3), let $v_{k-1,j}=F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$ and $u=F(\widetilde{y}_{k}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),y_{k-1}=0)$, then the numerator in (20) can be written as:
[…]
$=C(v_{k-1j},1)-C(v_{k-1j},q_{k-1,j})$
$=v_{k-1j}-C(v_{k-1j},q_{k-1,j})$
where the copula $C$ is $C_{L_{j}\widetilde{Y}_{k}\mid\bar{L}_{j-1},do(\bar{A}_{k}),Y_{k-1}=0}(\cdot;\Theta_{k-1,j})$, specifically. ∎
Appendix C Proof of Proposition 5.2
@@
Initialize the algorithm by specifying the cumulative distribution functions $F_{L_{0}}$, $F_{A_{0}\mid L_{0}}$, $F_{\widetilde{Y}_{1}\mid do(A_{0})}$, and the copula $C_{\widetilde{Y}_{1}L_{0}\mid do(A_{0})}$, parameterized by $\Theta_{0,0}$. Draw $w_{0,1}$, $w_{0,2}$, $w_{0,3}\sim U(0,1)$ and compute:
$l_{0}=F^{-1}_{L_{0}}(w_{0,1})$ $a_{0}=F^{-1}_{A_{0}\mid L_{0}}(w_{0,2}\,|\,l_{0}),$
@@
$F_{\widetilde{Y}_{1}\,|\,do(A_{0})}(\widetilde{y}_{1}\,|\,do(a_{0}))$ $=h^{-1}\big{\{}F_{\widetilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\widetilde{y}_{1}\,|\,l_{0},do(a_{0})),F_{L_{0}}(l_{0});\Theta_{0,0}\big{\}},$
set $w_{0,3}=F_{\widetilde{Y}_{1}\,|\,L_{0},do(A_{0})}(\widetilde{y}_{1}\,|\,l_{0},do(a_{0}))$ and substitute $F_{L_{0}}(l_{0})$ with $w_{0,1}$, yielding:
$u_{1}=h^{-1}\left(w_{0,3},w_{0,1};\Theta_{0,0}\right).$
By inversion, compute $\widetilde{y}_{1}=F_{Y_{1}\mid do(A_{0})}^{-1}(u_{1})$, and the corresponding binary outcome $y_{1}=I(\widetilde{y}_{1}<1)$. This completes the first iteration, yielding a sample $(l_{0},a_{0},\widetilde{y}_{1},y_{1})$. If $y_{1}=1$, stop; otherwise, before proceeding to $k=1$, we store the following values:
$u_{0,0}=F_{L_{0}}(l_{0})$ $q_{0,0}=F_{\widetilde{Y}_{1}\mid do(A_{0})}(\widetilde{y}_{1}=1\,|\,do(a_{0})).$
For $k=1$, define the necessary cumulative distribution functions $F_{L_{1}\,|\,L_{0},A_{0}}$, $F_{A_{1}\mid\bar{L}_{1},A_{0}}$, $F_{\widetilde{Y}_{2}\mid do(\bar{A}_{1})}$, and copulas $C_{\widetilde{Y}_{2}L_{1}\mid L_{0}do(\bar{A}_{1})}$ and $C_{\widetilde{Y}_{2}L_{0}\mid do(A_{0})}$, parameterized by $\Theta_{1,1}$ and $\Theta_{1,0}$, respectively. These two copulas associate $\widetilde{Y}_{2}$ with previous time-varying confounders. Draw $w_{1,1}$, $w_{1,2}$, $w_{1,3}$ from $U(0,1)$, and obtain $(l_{1},a_{1})$ as before.
We now need to obtain a sample of $\widetilde{y}_{2}$. Calculate:
$F(l_{0}\,|\,do(a_{0}),y_{1}=0)$ $=\frac{u_{0,0}-C^{*}_{\widetilde{Y}_{1}L_{0}\mid A_{0}}(u_{0,0},q_{0,0};\Theta_{0,0})}{1-q_{0,0}}$ (21)
$F(\widetilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\widetilde{y}_{2}\,|\,\bar{l}_{1},do(\bar{a}_{1}),y_{1}=0),F(l_{1}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0);\Theta_{1,1}\big{\}}$
$F(\widetilde{y}_{2}\,|\,do(a_{1}),y_{1}=0)$ $=h^{-1}\big{\{}F(\widetilde{y}_{2}\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0),F(l_{0}\,|\,do(\bar{a}_{1}),y_{1}=0);\Theta_{1,0}\big{\}}.$
Note that we omit some subscripts on quantiles for brevity. In (21) we derive the quantile of $L_{0}$ given $a_{0}$ among survivors the start of the $k=1$ interval, where $u_{0,0}$ and $q_{0,0}$ are already computed in step $k=0$. All other necessary quantities are specified at $k=1$. These computations yields a sample $(l_{1},a_{1},\widetilde{y}_{2},y_{2})$. If $y_{2}=1$, the simulation stops; otherwise, we store and compute the following values:
$u_{1,0}$ $=F_{L_{0}\mid Y_{1}}(l_{0}\,|\,y_{1}=0)$
$q_{1,0}$ $=F_{\widetilde{Y}_{2}\mid do(\bar{A}_{1})Y_{1}}(\widetilde{y}_{2}=1\,|\,do(\bar{a}_{1}),y_{1}=0)$
$u_{1,1}$ $=F_{L_{1}\mid L_{0}Y_{1}}(l_{0}\,|\,l_{0},y_{1}=0)$
$q_{1,1}$ $=F_{\widetilde{Y}_{2}\mid L_{0}do(\bar{A}_{1})Y_{1}}(\widetilde{y}_{2}=1\,|\,l_{0},do(\bar{a}_{1}),y_{1}=0)$
$=h\big{\{}q_{1,0};u_{1,0};\Theta_{1,0}\big{\}},$
where we simply store the first three quantities, which are directly available from the current step, while only the last one requires computation. Therefore, the base case is satisfied since the algorithm simulates the first time step and provides the required inputs for the next.
@@
[…]
$B\sim\operatorname{Bernoulli}(0.7)$ $L_{0}\sim\operatorname{Gamma}\!\left(k=(25+0.8B),\;\theta=2.5\right)$
$A\sim\operatorname{Bernoulli}(0.5)$ $S_{0}\sim\operatorname{Bernoulli}(0.1).$
The choice of these parameter values are loosely based on the baseline characteristics of SOUL reported in McGuire et al. (2023). For subsequent follow-up visits $k>0$:
$S_{k}$ $\sim\operatorname{Bernoulli}(\operatorname{expit}(\alpha_{0}+\alpha_{1}\,B+\alpha_{2}\,L_{k}+\alpha_{3}\,A))$
$L_{k}$ $\sim\operatorname{Gamma}(k=400(L_{k-1}+0.2\,B-0.5\,S_{k-1}-A_{k-1}),\;\theta=\frac{1}{400}).$
We target a Cox marginal structural model given by:
$\lambda(\widetilde{y}_{k+1}\,|\,do(A,\bar{S}_{k}),Y_{k}=0)=\lambda_{0}\exp\left(\beta_{A}\,A+\beta_{S}\,S_{k}+\beta_{AS}\,AS_{k}\right)=g_{k}(A,S_{k}),$
This parameterization assumes a constant hazard rate between follow-up visits, influenced by the uses of semaglutide and SGLT2-i at the most recent visit. This implies that given survival at follow-up time $k$, the incremental survival time $\widetilde{Y}_{k+1}$ follows an exponential distribution with the rate equal to the hazard above:
$\widetilde{Y}_{k+1}\sim\operatorname{Exp}\big{(}g_{k}(A,S_{k})\big{)}.$
If $\widetilde{Y}_{k+1}<1$ then $Y_{k+1}=1$, and 00 otherwise.
To understand the impact of differential use of SGLT2-i on the estimation of treatment effect of the primary treatment, semaglutide, we create several scenarios with different prevalence of SGLT2-i in the placebo arm. We achieve this by varying the coefficients ($\alpha_{0}$,$\alpha_{1}$,$\alpha_{2}$ and $\alpha_{3}$) of the model for $S_{k}$. We also explored varying the efficacy of SGLT2-i. Table 6 below describes 4 levels of prevalence and 3 levels of efficacy of SGLT2-i. In total, we simulate 10 scenarios.
SGLT2-i Prevalence Scenarios Baseline $S_{k}\sim\text{Bernoulli}(\text{expit}(-9.5+0.5B+0.1L_{k}-1.5A))$. Annual take-up: 10% in placebo, 2.5% in semaglutide group. High drop-in $S_{k}\sim\text{Bernoulli}(\text{expit}(-9+0.5B+0.1L_{k}-2A))$. Annual take-up: 15% in placebo, 2.5% in semaglutide group. Low drop-in $S_{k}\sim\text{Bernoulli}(\text{expit}(-10+0.5B+0.1L_{k}-A))$. Annual take-up: 5% in placebo, 2.5% in semaglutide group. No drop-in $S_{k}=0$ Effect Scenarios Moderate effect $\beta_{S}=\log(0.9)$, $\beta_{AS}=0.05$. Strong effect $\beta_{S}=\log(0.85)$, $\beta_{AS}=0.1$. Weak effect $\beta_{S}=\log(0.95)$, $\beta_{AS}=0.01$.
Table 6: Descriptions of SGLT2-i prevalence and efficacy scenarios. The prevalence scenarios vary the annual uptake of SGLT2-i in the placebo and semaglutide groups, while efficacy scenarios are defined by combinations of $\beta_{S}$ and $\beta_{AS}$, representing the isolated and interaction effects, respectively.
The values of $\beta_{S}$ are based on the findings of McGuire et al. (2021), who reported that SGLT2 inhibitors were associated with a reduced risk of MACE, with a hazard ratio of $0.90$ ($95\%$ confidence interval: 0.85–0.95) from a meta-analysis of six clinical trials. The interaction term represents the expectation that the combined effect of using both treatments is less than the total of their individual effects.
For each simulated dataset, we fit the above Cox marginal structural model $\lambda(\widetilde{y}_{k+1}\,|\,do(A,\bar{S}_{k}),Y_{k}=0)=\lambda_{0}\exp\left(\beta_{A}\,A+\beta_{S}\,S_{k}+\beta_{AS}\,AS_{k}\right)$ using stabilized inverse-probability-of-treatment weights (Hernán et al., 2000), calculated as
$W_{k}=\prod_{j=0}^{k}\frac{\hat{p}(S_{j}\mid A,\bar{S}_{j-1},Y_{j}=0)}{\hat{p}(S_{j}\mid B,A,\bar{L}_{j},\bar{S}_{j-1},Y_{j}=0)}.$ (22)
Remark D.1.
The denominator in the weight formula represents the correctly specified propensity score model under the assumed data-generating process, while the numerator stabilizes the weights. To avoid unintended confounding, the baseline covariates included in the numerator must also be adjusted for in the marginal structural model (Cole and Hernán, 2008). In accordance with the model in (23), we only include $A$ in the numerator for stabilization.
Appendix E Seaman and Keogh (2024) simulations
We present an example that broadly replicates the setting described in Seaman and Keogh (2024). We consider baseline variables $B$ and $C$, two time-varying covariates $L_{k}$ and $W_{k}$, and treatment $A_{k}$ at follow-up $k$. The time-to-event is $T$, with $Y_{k+1}=I(T<k+1)$ indicating failure status observed at the end of the $k$th interval. The auxiliary continuous variable $\widetilde{Y}_{k+1}$ is defined as previously. Specifically we choose that at baseline: